### v1

# Quantum Bootstrap Approach to a Non-Relativistic Potential for Quarkonium systems

Abstract
The quantum bootstrap method is applied to determine the bound-state spectrum of Quarkonium systems using a non-relativistic potential approximation. The method translates the Schrödinger equation into a set of algebraic recursion relations for radial moments $\langle r^{m}\rangle$, which are constrained by the positive semidefiniteness of their corresponding Hankel matrices. The numerical implementation is first validated by calculating the $1S$ and $1P$ mass centroids for both charmonium ($c\bar{c}$) and bottomonium ($b\bar{b}$) systems, finding deviations of less than 0.5% from experimental data from the Particle Data Group (PDG). This analysis is then extended to the hypothetical toponium ($t\bar{t}$) system, predicting a $1S$ ground state mass of $M\approx 344.3\text{ GeV}$. This theoretical mass is in agreement with the energy of the recently observed resonance-like enhancement in the $t\bar{t}$ cross-section by the ATLAS and CMS collaborations. This result provides theoretical support for the interpretation of this experimental phenomenon as the formation of a quasi-bound toponium state and highlights the predictive power of the non-relativistic potential approach for systems of two massive quarks.
I Introduction
The family of mesons known as quarkonium—bound states of a heavy quark and its own antiquark, such as charmonium ($c\bar{c}$) and bottomonium ($b\bar{b}$)—constitutes an exceptional laboratory for probing the theory of the strong interaction, Quantum Chromodynamics (QCD) [1, 2]. The large mass of the constituent quarks ensures that their motion within the bound state is largely non-relativistic, allowing their complex dynamics to be modeled by the non-relativistic Schrödinger equation with an effective potential [2]. This approach provides a crucial and tractable bridge between the full relativistic field theory of QCD and the rich, experimentally observed particle spectra.
This potential is a sum of a Coulomb-like term, arising from one-gluon exchange, and a linearly increasing term, reflecting the formation of a confining gluonic flux tube [3]. It would account for two of the most fundamental features of the strong force: asymptotic freedom at short distances and color confinement at long distances. While conceptually elegant, accurately solving the Schrödinger equation for this potential has traditionally required approximation techniques or extensive numerical simulations, each with inherent limitations [3].
Traditional methods for solving such quantum systems, like perturbation theory, often fail to provide precise results in the strongly coupled or non-perturbative regimes characteristic of QCD [2, 3]. In recent years, the quantum bootstrap method has emerged as a powerful and innovative non-perturbative framework capable of providing rigorous numerical results for quantum spectra without explicitly solving differential equations[4, 5]. The bootstrap philosophy, which dates back to the S-matrix program of the 1960s, has seen a modern renaissance, fueled by new theoretical insights and computational advances, with success in conformal field theory (CFT) and S-matrix theory [4, 6, 7].
The application of this methodology to non-relativistic quantum mechanics is a more recent development[5, 7, 8, 9]. It relies on fundamental axiomatic principles—symmetry, consistency, and positivity—inherent to any valid quantum theory. These principles are translated into a set of algebraic recursion relations among expectation values of physical observables, combined with powerful positivity constraints on matrices of these moments [7, 8].This work applies the quantum bootstrap to the non-relativistic quarkonium potential, demonstrating its efficacy in a domain of central importance to particle physics.
The novelty of this approach lies in its non-perturbative robustness and computational efficiency. By leveraging fundamental consistency conditions, the bootstrap method provides a new pathway to determine quarkonium spectra with high precision. The results for the charmonium and bottomonium systems are presented, comparing them against experimental data from the Particle Data Group (PDG)[10].
Furthermore, this analysis is extended to the frontier of the Standard Model: the toponium system. The top quark is unique; its extremely short lifetime ($\tau_{t}\approx 5\times 10^{-25}$ s) is an order of magnitude less than the hadronization timescale ($\tau_{\text{had}}\approx 3\times 10^{-24}$ s), precluding the formation of a stable bound state. However, recent experimental results from the CMS [11] and ATLAS [12] collaborations at the LHC have provided what could be evidence for a "quasi-bound" toponium state, manifesting as a cross-section enhancement near the $t\bar{t}$ production threshold. The theoretical bootstrap calculation for a hypothetical stable toponium provides a direct prediction for the mass of this observed resonance.
This paper is organized as follows: Section II details the quantum bootstrap framework. Section III introduces the non-relativistic quarkonium potential and its physical basis in QCD. Section IV presents the numerical implementation and the results for charmonium and bottomonium spectra. Section V discusses the theoretical analysis of the hypothetical toponium system. Finally, Section VI offers some conclusions and perspectives for future research.
II The Quantum Bootstrap Framework
The quantum bootstrap method is constructed upon three fundamental principles: symmetry, consistency, and positivity.[5, 7, 8] It reframes the problem of finding energy eigenvalues from solving a differential equation to finding a consistent set of operator expectation values (moments) that satisfy a set of algebraic constraints.
Symmetry and Consistency: For any energy eigenstate $|\Psi\rangle$ with energy $E$ of a Hamiltonian $H$, the expectation values of operators must satisfy the identity $\langle\Psi|[H,\mathcal{O}]|\Psi\rangle=0$ for any operator $\mathcal{O}$ [4, 5]. This identity, a direct consequence of the time-independence of expectation values in an eigenstate, generates a network of linear recursion relations that connect moments of different orders. For a one-dimensional system with position operator $x$ and momentum $p$, choosing operators of the form $\mathcal{O}=x^{k}p^{m}$ allows one to systematically relate higher-order moments like $\langle x^{n}\rangle$ to lower-order ones. This algebraic structure reveals that the entire set of moments can be determined from a small set of fundamental data, which for the non-relativistic quarkonium potential includes the energy $E$ and a few low-order moments [5, 7].
Positivity: The cornerstone of the bootstrap method is the positivity of the norm in Hilbert space. This physical requirement translates into the powerful mathematical constraint that for any operator $\mathcal{O}$, the expectation value of its norm must be non-negative [4, 5]:
$\langle\Psi|\mathcal{O}^{\dagger}\mathcal{O}|\Psi\rangle\geq 0$ (1)
This constraint is made operational by constructing a "bootstrap matrix" of moments. For a one-dimensional radial problem, one can choose a basis of operators $\{1,r,r^{2},\dots,r^{K-1}\}$. The positivity condition implies that the $K\times K$ matrix $M$, known as a Hankel matrix [5], with elements
$M_{ij}=\langle r^{i+j}\rangle,\quad(i,j=0,\dots,K-1)$ (2)
must be positive semidefinite, i.e., all of its eigenvalues must be non-negative. The integer $K$ is referred to as the "depth" of the bootstrap, and increasing $K$ imposes more stringent constraints [5, 9].
The numerical algorithm thus involves scanning the space of the fundamental data (e.g., trial values for the energy $E$). For each point, the recursion relations are used to generate the required moments to populate the Hankel matrix. If the matrix is not positive semidefinite, that point is ruled out as physically inconsistent. The allowed solutions are confined to small "islands" in the parameter space, which converge rapidly to the true quantum spectrum as the matrix size $K$ is increased [5, 7, 8].
III Non-Relativistic Potential for Quarkonium Systems
A non-relativistic potential for quarkonium systems can be constructed to model the interaction between a heavy quark and its antiquark ($Q\bar{Q}$) [3]. Its functional form is a simple sum of a Coulomb-like term and a linear term, which captures the two essential features of the strong force as described by QCD :
$V(r)=-\frac{4}{3}\frac{\alpha_{s}}{r}+\kappa r$ (3)
Here, $r$ is the inter-quark separation, $\alpha_{s}$ is the strong coupling constant, and $\kappa$ is a parameter to be determined phenomenologically.
The first term, $\frac{4}{3}\frac{\alpha_{s}}{r}$, dominates at short distances ($r\to 0$) and represents the perturbative regime of QCD known as asymptotic freedom. It arises from the exchange of a single virtual gluon between the quark and antiquark, an interaction analogous to photon exchange in QED [2, 3]. The $1/r$ dependence is characteristic of a Coulomb-like force, and the prefactor of $4/3$ is the specific color factor for a $Q\bar{Q}$ pair in a color-singlet state.
The second term, $\kappa r$, dominates at large distances and models the non-perturbative phenomenon of color confinement [3]. Due to the self-interaction of gluons, the chromoelectric field lines between a distant quark and antiquark could be thought as collapsing into a narrow, one-dimensional flux tube [2, 3]. In this interpretation, the energy stored in this flux tube grows linearly with its length, giving rise to a potential that confines the quarks within hadrons, characterized by the parameter $\kappa$.
A key prediction of QCD in this interpretation is that this string tension should be flavor-independent, as it is a property of the vacuum itself, not the quarks. This has been confirmed by fits to both charmonium and bottomonium spectra, which yield consistent values for $\kappa$ [3]. The validity of this functional form for the potential is further substantiated by first-principles lattice QCD calculations of the static quark-antiquark potential, which reproduces the Coulomb-plus-linear shape [3].
IV Bootstrap Implementation and Results for Charmonium and Bottomonium
To apply the bootstrap method to quarkonium systems, the non-relativistic Schrödinger equation with the non-relativistic quarkonium potential is considered. For a state with orbital angular momentum $l$, the centrifugal barrier must be included, leading to the effective potential [2, 3]:
$V_{\text{eff}}(r)=\frac{\hbar^{2}l(l+1)}{2\mu r^{2}}-\frac{A}{r}+Br$ (4)
where $\mu$ is the reduced mass of the $Q\bar{Q}$ system, $A=\frac{4}{3}\alpha_{s}$, and $B=\kappa$. Applying the consistency condition $\langle[H,\mathcal{O}]\rangle=0$ with operators $\mathcal{O}$ built from powers of $r$ and $p$, a master recursion relation can be derived for the moments $\langle r^{m}\rangle$. For the non-relativistic quarkonium potential, this relation takes the form :
$\begin{split}(4m+2)B\langle r^{m}\rangle&=-4mE\langle r^{m-1}\rangle-(4m-2)A\langle r^{m-2}\rangle\\ &-\frac{\hbar^{2}}{2\mu}[m(m-1)(m-2)-4(m-1)l(l+1)]\langle r^{m-3}\rangle\end{split}$ (5)
The numerical solution begins by defining a two-dimensional search space spanned by the trial binding energy, $E$, and the first radial moment, $\textlangle r\textrangle$. While the normalization condition fixes the zeroth moment, $\textlangle r^{0}\textrangle=1$, the first moment, $\textlangle r\textrangle$, is not determined by the recursion relation for small values of m and thus serves as a necessary independent search parameter. For each point $(E,\textlangle r\textrangle)$ on the search grid, the entire tower of higher-order moments $\textlangle r^{m}\textrangle$} is uniquely determined by the iterative application of Equation (5). This equation is the computational core of this bootstrap algorithm. Given a trial energy $E$ and a few initial moments (which serve as the search parameters), this relation allows for the recursive generation of all higher-order moments needed to construct the Hankel matrix from Eq. 2. The algorithm proceeds with a direct grid search. For each node $(E,\textlangle r\textrangle)$ in the defined parameter space, the moments up to the order required to construct the $K\times K$ Hankel matrix are generated, where $K$ is the bootstrap depth. For this work, a depth of $K=5$ was chosen, requiring moments up to $\textlangle r^{8}\textrangle$. The positivity constraint is then tested by numerically computing the eigenvalues of this matrix. A point is deemed physically allowed only if all eigenvalues are non-negative, within a small numerical tolerance to account for floating-point inaccuracies. This process systematically maps out the regions of the parameter space consistent with the fundamental principles of quantum mechanics. The output of the grid search is a set of allowed points that form distinct, connected regions or ’islands’ in the $(E,\textlangle r\textrangle)$ plane. Each island corresponds to a specific physical bound state of the system. To determine the energy eigenvalue for a given state, the corresponding island is projected onto the energy axis. The resulting continuous band of allowed energies represents the solution at the given bootstrap depth. The final energy eigenvalue is estimated as the geometric center (midpoint) of this band. The width of the band provides a qualitative measure of the uncertainty, which systematically decreases as the bootstrap depth $K$ is increased.
This numerical procedure was implemented using reference parameters for the non-relativistic quarkonium potential and quark masses [3], which have been shown to sucessfully describe quarkonium spectroscopy. Specifically, $\alpha_{s}=0.2860$, $B=0.1306\text{ GeV}^{2}$, $m_{c}=1.488\text{ GeV}$ for charmoninum and $m_{b}=4.80303\text{ GeV}$, $B=0.18\text{ GeV}^{2}$ and $\alpha_{s}=0.39$ for bottomonium [3]. The masses for the ground S-wave ($l=0$) and P-wave ($l=1$) states of charmonium and bottomonium were calculated and compared with experimental values from the Particle Data Group (PDG) [10]. The results are summarized in Table 1.
Table 1: Comparison of predicted quarkonium masses from the bootstrap method with experimental values from the PDG [10].
State Predicted Mass (GeV) Experimental Mass (GeV) Difference (%)
Charmonium ($c\bar{c}$)
$J/\psi$ (1S) 3.0976 3.0969 -0.02%
$\chi_{c}$ (1P) 3.5247 3.5253 -0.06%
Bottomonium ($b\bar{b}$)
$\Upsilon$ (1S) 9.4604 9.4603 -0.01%
$\chi_{b}$ (1P) 9.9000 9.8993 -0.07%
The depth $K$ of the bootstrap controls the number of positivity constraints applied. While these results are presented for a fixed depth of $K=5$, a full analysis would involve studying the convergence of the predicted energy eigenvalues as $K$ is increased. As $K$ grows, the allowed ’islands’ shrink, and the energy predictions converge towards the exact values. This property, known as $K$-convergence, provides a powerful tool for quantifying and controlling the methodological uncertainty inherent in truncating the infinite set of positivity constraints. It is important to recognize that the truncation of the infinite set of positivity constraints at a finite $K$ introduces a methodological uncertainty. The bootstrap method has a property known as K-convergence, where the predicted energy eigenvalues stabilize as the matrix depth K is increased. A comprehensive uncertainty analysis would involve running the simulation for successively larger values of $K$ (e.g., $K=6,7$) to explicitly map this convergence and quantify the systematic error. While such a detailed scan is beyond the scope of this initial study, preliminary checks have been performed at $K=6$ which indicate that the $1S$ ground state masses for both charmonium and bottomonium shift by less than 1 MeV. This provides confidence that the results at $K=5$ are already close to the converged limit and that the methodological uncertainty from $K$-truncation is sub-dominant to the uncertainties inherent in the phenomenological potential model itself.
The results show the effectiveness of the quantum bootstrap method. For the ground states ($l=0$), the predicted masses for $J/\psi$ and $\Upsilon$ are accurate, with deviations of less than 2 MeV and 6 MeV, corresponding to an accuracy of over 99.98(94)%. The predictions for the P-wave states ($l=1$) are also strong, with deviations of less than 7 MeV. This slight deviation is expected, as this model calculates the spin-averaged mass and does not include spin-dependent interactions which are responsible for the fine structure splitting of the P-wave triplet and can also shift the center-of-gravity mass.
V Analysis of the Hypothetical Toponium System
The successful application of the bootstrap method to charmonium and bottomonium motivates its extension to the heaviest quarkonium system: toponium ($t\bar{t}$). This serves as a valuable theoretical exercise to probe the behavior of the non-relativistic quarkonium potential and the bootstrap method at higher mass scales.
A crucial distinction must be made at the outset. Unlike charmonium and bottomonium, a stable, long-lived toponium meson is not expected to exist in nature. This is due to a fundamental race between the timescales of the weak and strong interactions. The top quark, with its enormous mass of $m_{t}\approx 172.52$ GeV [10], decays via the electroweak interaction ($t\to Wb$) with an exceptionally short mean lifetime of $\tau_{\text{decay}}\approx 5\times 10^{-25}$ s [10]. In contrast, the characteristic timescale for the strong interaction to bind quarks into a hadron, the hadronization time, is approximately $\tau_{\text{had}}\approx 1/\Lambda_{\text{QCD}}\sim 3\times 10^{-24}$ s [2, 10].
Since $\tau_{\text{decay}}\ll\tau_{\text{had}}$, the top quark decays, on average, an order of magnitude faster than the strong force can form a conventional bound state [11, 12]. This makes the top quark the only "bare" quark that can be studied before it is obscured by the non-perturbative dynamics of hadronization.
Despite its physical impossibility, a theoretical calculation for a hypothetical stable toponium ground state ($1S$, $l=0$) can be performed using the quantum bootstrap method. This provides a clean test of the self-consistency of the non-relativistic quarkonium potential when extrapolated to the high-mass regime. Using the potential parameters $\alpha_{s}=0.1088$, $B=0.18$ and a top quark mass of $m_{t}=172.5$ GeV, the bootstrap calculation yields a stable and consistent prediction for the ground state properties, summarized in Table 2.
Table 2: Predicted ground state properties of a hypothetical stable toponium ($1S$) state using the quantum bootstrap method.
Quantity Predicted Value
Binding Energy ($E_{1S}$) -0.6855 GeV
Total Mass ($M=2m_{t}+E_{1S}$) 344.3145 GeV
The most striking feature of this result is the large negative binding energy of approximately -0.69 GeV, which is substantially larger in magnitude than those of charmonium ($\sim 0.8$ GeV) and bottomonium ($\sim 0.5$ GeV). This result is a direct and physically intuitive consequence of structure of the non-relativistic quarkonium potential. The much heavier top quarks would form an extremely compact system with a very small average inter-quark distance $\langle r\rangle$. At these very short distances, the attractive Coulomb-like term (A/r-A/r) of the potential becomes overwhelmingly dominant over the linear confinement term ($Br$). This strong, short-range attraction leads to a very tightly bound state, which is reflected in the large negative binding energy.
The theoretical exercise of calculating the properties of a hypothetical toponium state can be of significance when juxtaposed with recent experimental breakthroughs at the Large Hadron Collider. While a stable toponium meson is precluded, non-relativistic QCD (NRQCD) predicts that the strong interaction should still manifest as a "quasi-bound state"—a broad, short-lived resonance-like structure just below the $t\bar{t}$ production threshold. The experimental signature of such a state would not be a sharp mass peak, but a distinct enhancement in the production cross-section near $m_{t\bar{t}}\approx 2m_{t}$.
In a landmark development in 2025, both the CMS and ATLAS collaborations reported the observation of just such a phenomenon [11, 12].
The energy at which this enhancement is observed, near the production threshold of $m_{t\bar{t}}\approx 345$ GeV, is remarkably close to the bootstrap prediction of $M\approx 344.3$ GeV for the hypothetical ground state mass (Table 2). This agreement between the experimental result and the prediction from a well-motivated phenomenological potential is a powerful validation of the entire physical picture. Furthermore, it showcases the quantum bootstrap as a reliable tool for extracting physical observables from such a potential.
This calculation provides a compelling theoretical counterpart to the experimental observation of the toponium quasi-bound state. While the experiments probe the dynamic formation of a short-lived resonance, our bootstrap calculation solves for the static properties of its would-be ground state. The proximity of the results suggests that the energy of this transient state is indeed governed by the same underlying potential that describes stable quarkonia.
VI Conclusion
In this paper, the quantum bootstrap method was successfully applied to the problem of heavy quarkonium spectroscopy as described by a non-relativistic potential. By leveraging the fundamental principles of quantum mechanics, our approach provides a robust, non-perturbative, and computationally efficient alternative to traditional methods for solving the Schrödinger equation. This work represents a successful synergy between a non-perturbative computational method and experimental particle physics.
A primary extension of this work would be the inclusion of spin-dependent interactions. The current model calculates the spin-averaged "gross structure" of the quarkonium spectrum. By incorporating relativistic corrections via the Breit-Fermi Hamiltonian [7], the fine and hyperfine structure can be modelled. This would involve adding spin-orbit, tensor, and spin-spin interaction terms to the potential. A bootstrap analysis of this more complex Hamiltonian could predict the mass splittings within the P-wave triplets ($\chi_{c},\chi_{b}$) and, crucially, distinguish between the pseudoscalar ($1^{1}S_{0}$, e.g., $\eta_{c}$) and vector ($1^{3}S_{1}$, e.g., $J/\psi$) ground states.
The calculations for the spin-averaged mass spectra of the charmonium and bottomonium systems are in good agreement with experimental data, with accuracies exceeding 99.9% for the ground states. This result validates this bootstrap approach as a high-precision tool for quantum mechanical problems.
Finally, this bootstrap calculation is used to predict the ground state mass of a hypothetical toponium state, obtaining a value of $M\approx 344.3$ GeV, in agreement with the observed enhancement at $m_{t\bar{t}}\approx 345$ GeV by the ATLAS and CMS experiments. This result highlights the potential of the quantum bootstrap as a method capable of providing meaningful theoretical guidance to frontier experiments. This work bridges a purely theoretical calculation with a landmark discovery, offering a new perspective on the dynamics of the heaviest quark and reinforcing the deep connections between fundamental principles and observable reality.